OpenStudy (flygirlmandi):

x+19 is less than or equal to -5 ? what is x a=x is less than or equal to -24 b=x is less than or equal to 24 c=x is less than or equal to 14 d=x is less than or equal to -14 thanks :)

OpenStudy (lynfran):

\[x+9\le-5\]

OpenStudy (lynfran):

to find x we have to carry the 9 to the the right side of the inequality by ?

OpenStudy (lynfran):

subtracting 9 from both sides

OpenStudy (flygirlmandi):

so 9 away from -5 ? and then 9 away from 19 ? im really bad at this stuff, sorry :)

OpenStudy (lynfran):

o i see .. i made an error didnt see 19 lol saw the 9 alone... so it shd be \[x+19 \le-5\]

OpenStudy (lynfran):

we need to subtract 19 from both sides to find x

OpenStudy (flygirlmandi):

so 19 away from -5 and 19 away from 19 ?

OpenStudy (lynfran):

that way the 19 from the left side will go to zero leaving the "x" term alone

OpenStudy (lynfran):

yep

OpenStudy (lynfran):

so we have \[x+19-19\le-5-19\]

OpenStudy (lynfran):

19-19 will give 0 whats -5-19?